OpenStudy (anonymous):

George deposited $1,300.00 in a savings account with an interest rate of 5% for 6 years. The interest is compounded daily. How much money did George have at the end of the six years? A. $1,754.56 B. $1,754.78 C. $6,571.13 D. $6,556.21

OpenStudy (anonymous):

\[1,300\times \left(1+\frac{.05}{365}\right)^{365\times 6}\] and a calculator

OpenStudy (anonymous):

i dont have one with me

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

you are on a computer so of course you have a calculator http://www.wolframalpha.com/input/?i=1%2C300%5Ctimes+%5Cleft%281%2B%5Cfrac%7B.05%7D%7B365%7D%5Cright%29%5E%7B365%5Ctimes+6%7D

OpenStudy (anonymous):

ohh lol